leilanithomas2024:

What is the equation of the line that has a slope of 3 and goes through the point (-3,-5)? A. y = 3x + 4 B. y = 3x − 14 C. y = 3x − 4 D. y = 3x + 12

mxddi3:

so we do this with the point-slope formula of \[y-y1=m(x-x1)\] m, the slope, in this case is -3 and we have (-3,-5) we label the ordered pair as x1 and y1. do you follow so far?

leilanithomas2024:

Yes

leilanithomas2024:

Because the formula to a slope is y2-y1/x2-x1 am i right? @mxddi3

mxddi3:

ok good. now, since it's y-y1=m(x-x1 and we are given the slope of 3, our new equation is : \[y-(-5)=3(x- [-3]) \] ok, now, can you simplify this or do you need help?